OpenStudy (anonymous):

Solve the equation: x^2+3x=0 (show steps)

OpenStudy (anonymous):

\[x ^{2}+3x=x(x+3)=0\]Then, you have two things that when multiplied together make zero, so, one of them must be zero. So, either:\[x=0\]or\[x+3=0=>x=-3\]

OpenStudy (anonymous):

what is the command for the arrow?

OpenStudy (anonymous):

so its 0 and -3

OpenStudy (anonymous):

I just use an equals and then a greater than sign.

OpenStudy (anonymous):

Yes, 0 OR -3.